# Quantum Kernel Methods under Scrutiny: A Benchmarking Study

### v1

@@ D.2 Details on data encoding strategies
Figure A15 provides insights into the effect of the two different encoding strategies for the hidden manifold diff dataset with $m=13$. The findings discussed for the two curves diff dataset in Sec. IV.3 still apply.
@@
D.3 Details on in-depth analysis of PQK design options
@@
D.3.1 Overview on test performances
@@
Figure A17: ROC-AUC test score as a function of the measurement operators given in Eqs. (14),  (16) and (18) for Gaussian, Matér and RationalQuadratic outer kernel functions $\kappa$. (a) two curves diff dataset for $d=4$ and degree $D=13$. (b) hidden manifold diff dataset for $d=4$ and manifold dimension $m=13$. The data for each $O^{k}$ are aggregated across the encoding circuits considered in this work.
@@
D.3.2 Extended correlation analysis
In order to support the efforts in Sec. IV.4 to gain a deeper understanding of what are basic mechanisms in PQKs responsible for learning, we provide all (statistically significant) semi-partial correlations of the various hyerparameters of the QSVR/QSVC-PQK model for a given outer kernel $\kappa$, both among themselves and with model test performance in Tab. 2. Here, we note that while negative correlation coefficients w.r.t. the MSE point at overall better model performance, it is the other way round for correlations with ROC-AUC scores.
Table 2: Statistically significant semi-partial correlations of the various hyperparameter of the QSVR/QSVC-PQK model for a given outer kernel function $\kappa$, both among themnselves and with test performance scores. This table supports the corresponding discussion in Ref. IV.4. We note that while negative correlation coefficients w.r.t. the MSE point at overall better model performance, it is the other way round for correlations with ROC-AUC scores.
@@ D.3.2 Extended correlation analysis
D.3.3 Complementary plots to Figure 12
In Sec. IV.4 we thoroughly investigate the underlying mechanisms that are eventually responsible for learning in PQK approaches. While comprehensive correlation analyses in Tabs. 1 and 2 provide profound insights, a deeper understanding of these findings follows from studying the distance measure as defined in Eq. (19). For Gaussian outer kernels we can additionally define the matrix of the projected quantum circuit contributions $F_{\bm{\theta}}$ as given in Eq. (20). In the main text, we detail the corresponding discussion for the QFMNIST ($n=8$) dataset in case of the Gaussian outer kernel. The central illustration for this is provided in Fig. 12. Subsequent to these explanations we display the corresponding findings for the Matérn and RationalQuadratic outer kernels in Figs. A18 and A19, respectively.
Figure A18: Analogous illustration to Fig. 12 as discussed in Sec. IV.4 for the QFMNIST dataset with $n=8$ components and the Matérn outer kernel.
Figure A19: Analogous illustration to Fig. 12 as discussed in Sec. IV.4 for the QFMNIST dataset with $n=8$ components and the RationalQuadratic outer kernel.
@@ D.3.3 Complementary plots to Figure 12
Similarly, we explore all remaining datasets of this study and provide corresponding plots as follows:
  • Friedman with Gaussian, Matérn and RationalQuadratic outer kernels in Figs. A20A21 and A22
[…]
    NH3-PES with Gaussian, Matérn and RationalQuadratic outer kernels in Figs. A23A24 and A25
  • two curves diff with Gaussian, Matérn and RationalQuadratic outer kernels in Figs. A26A27 and A28
  • hidden manifold diff with Gaussian, Matérn and RationalQuadratic outer kernels in Figs. A29A30, and A31
Figure A20: Analogous illustration to Fig. 12 as discussed in Sec. IV.4 for the Friedman dataset with $n=10$ features and the Gaussian outer kernel.
Figure A21: Analogous illustration to Fig. 12 as discussed in Sec. IV.4 for the Friedman dataset with $n=10$ features and the Matérn outer kernel.
Figure A22: Analogous illustration to Fig. 12 as discussed in Sec. IV.4 for the Friedman dataset with $n=10$ features and the RationalQuadratic outer kernel.
[…]
Figure A26: Analogous illustration to Fig. 12 as discussed in Sec. IV.4 for the two curves diff dataset with degree $D=13$ and the Gaussian outer kernel.
Figure A27: Analogous illustration to Fig. 12 as discussed in Sec. IV.4 for the two curves diff dataset with degree $D=13$ and the Matérn outer kernel.
Figure A28: Analogous illustration to Fig. 12 as discussed in Sec. IV.4 for the two curves diff dataset with degree $D=13$ and the RationalQuadratic outer kernel.
Figure A29: Analogous illustration to Fig. 12 as discussed in Sec. IV.4 for the hidden manifold diff dataset with manifold dimension $m=13$ and Gaussian outer kernel.
Figure A30: Analogous illustration to Fig. 12 as discussed in Sec. IV.4 for the hidden manifold diff dataset with manifold dimension $m=13$ and Matérn outer kernel.
Figure A31: Analogous illustration to Fig. 12 as discussed in Sec. IV.4 for the hidden manifold diff dataset with manifold dimension $m=13$ and RationalQuadratic outer kernel.
@@ Appendix E Optimization of trainable parameters in data encoding circuits
The results listed in Tab. 3 indicate that optimization can lead to a significant improvement in the case of the two curves diff dataset, but, this improvement is not guaranteed, as demonstrated by the counterexample of the hidden manifold diff dataset.
@@
This discrepancy may be attributed to a poorly tuned optimizer. A detailed analysis of this issue, as well as the investigation of a broader range of datasets and data encoding circuits, however, is beyond the scope of this work. Nevertheless, we highlight the potential of KTA optimization for datasets with high complexity with correlating sub-optimal QKMs.

### v2

@@ D.4 Details on the Influence of Data Encoding Strategies
Figure A15 provides insights into the effect of the two different encoding strategies for the hidden manifold diff dataset with $m=13$. The findings discussed for the two curves diff dataset in Sec. IV.4 still apply.
@@
D.5 Details on the Analysis of PQK Design Options
@@
D.5.1 Overview on Test Performances
@@
Figure A17: ROC-AUC test score for QSVC-PQK models as a function of the measurement operators given in Eqs. (14),  (16) and (18) for Gaussian, Matér and RationalQuadratic outer kernel functions $\kappa$. This supports the discussion around Tab. 1 in Sec. IV.5 of the main text. The results corresponding to the following classification datasets are shown: (a) two curves diff dataset for $d=4$ and degree $D=13$. (b) hidden manifold diff dataset for $d=4$ and manifold dimension $m=13$. The data for each $O^{k}$ are aggregated across the encoding circuits considered in this work.
@@
D.5.2 Extended Correlation Analysis
In order to support the efforts in Sec. IV.5 to gain a deeper understanding of what are basic mechanisms in PQKs responsible for learning, we provide all (statistically significant) semi-partial correlations of the various hyerparameters of the QSVR/QSVC-PQK model for a given outer kernel $\kappa$, both among themselves and with model test performance in Tab. 2. Here, we note that while negative correlation coefficients w.r.t. the MSE point at overall better model performance, it is the other way round for correlations with ROC-AUC scores.
Table 2: Statistically significant semi-partial correlations of the various hyperparameter of the QSVR/QSVC-PQK model for a given outer kernel function $\kappa$, both among themnselves and with test performance scores. This table supports the corresponding discussion in Ref. IV.5. We note that while negative correlation coefficients w.r.t. the MSE point at overall better model performance, it is the other way round for correlations with ROC-AUC scores.
@@ D.5.2 Extended Correlation Analysis
D.5.3 Complementary Plots to Figure 12
In Sec. IV.5 we thoroughly investigate the underlying mechanisms that are eventually responsible for learning in PQK approaches. While comprehensive correlation analyses in Tabs. 1 and 2 provide profound insights, a deeper understanding of these findings follows from studying the distance measure as defined in Eq. (19). For Gaussian outer kernels we can additionally define the projected quantum circuit contributions $F_{\bm{\theta}}$ as given in Eq. (20). In the main text, we detail the corresponding discussion for the QFMNIST ($d=8$) dataset in case of the Gaussian outer kernel. The central illustration for this is provided in Fig. 12. In addition to these explanations we display the corresponding findings for the Matérn and RationalQuadratic outer kernels in Figs. A18 and A19, respectively.
Figure A18: Analogous illustration to Fig. 12 as discussed in Sec. IV.5 for the QFMNIST dataset with $d=8$ components and the Matérn outer kernel.
Figure A19: Analogous illustration to Fig. 12 as discussed in Sec. IV.5 for the QFMNIST dataset with $d=8$ components and the RationalQuadratic outer kernel.
@@ D.5.3 Complementary Plots to Figure 12
Similarly, we can provide corresponding discussions and illustrations for all remaining datasets of this study upon reasonable request.
@@ Appendix E Optimization of Trainable Parameters in Data Encoding Circuits
The results listed in Tab. 3 indicate that optimization can lead to a significant improvement in the case of the two curves diff dataset, but, this improvement is not guaranteed, as demonstrated by the counterexample of the hidden manifold diff dataset. This discrepancy may be attributed to a poorly tuned optimizer. A detailed analysis of this issue, as well as the investigation of a broader range of datasets and data encoding circuits, however, is beyond the scope of this work. Nevertheless, we highlight the potential of KTA optimization for datasets with high complexity with correlating sub-optimal QKMs.